North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Home Locomotive Histories Q6 63395 (NER T2 2238) Q6 Blog Saturday 16 January 2016 and Wednesday 13 January 2016

Saturday 16 January 2016 and Wednesday 13 January 2016

On Wednesday two old fresh faces joined the team in Mike Oliver and Brian Armstrong - a warm welcome back - with Les Harper, Jon Bradley, Derek Shorten, and Ian Pearson. Mike cleaned the right hand piston valve, and Brian cleaned the left hand one: both now done. Jon worked on cleaning and clearing the rest of the concrete from the smoke box - another job completed. Les removed the taper pin from the right hand trailing side rod, but then it was realised that the loco would have to be moved to get the other pins out, plus there was packing under the left hand slide bars, so that job was abandoned. Les then removed all the cylinder cock linkages and these are now on the bench in the workshop all marked up. The valves are still to remove.

Derek carried on cleaning the outside frames and also made a start underneath. Ian got the short straw again and continued cleaning the firebox. All of the front grate had to be removed with Brian's help, and these bars are now stacked in the tender. A large build up of ash had accumulated behind the grate bearer and foundation ring, and this has been removed to the ash pan.

There is still ash in the two blocked flue tubes although some more blowing out with compressed air was tried towards the end of the day. No water was available to wash these flues as it was turned off at the shed and the main tank on the tunnel top was empty.

The pit pump worked fine all day, but the angle grinder was not found and the valves were cleaned manually. New oxygen cylinders were delivered to the Railway about 2pm, but one has still to be moved to Deviation Shed.

For today (Saturday), Neal, James, Peter and Andy Lowes all made it to Grosmont in very slippery conditions. The commitment shown by those present in getting there was very much appreciated, especially as Andy had left his keys for the workshop at home! A big thank you to Andrew, the duty fitter, for allowing us to use some of the MPD gear. The pit was cleaned out by Peter who then assisted with stripping the cross head.

A new bottle of oxygen was obtained and put on our bottle stand in Deviation Shed, along with some new strops and shackles. Neal, James and Andy removed the left piston that had been released last week from the cross head and stood it up safely next to the locomotive. A start was then made on removing the one on the right side - this will be completed next week. A dimensional survey was done on both cylinders and slide bars so a plan can be put in place to continue the winter maintenance to a speedy conclusion.

There will be another working party on Wednesday, for which the jobs are:

  • Continue with cab valve refurbishment (Mike Oliver?).
  • Continue with cylinder drain cocks check fit to cylinders/steam chest and refurbish (Les Harper?).
  • Continue with side rod bush clearances using feeler gauges and record on side of rod (the taper pins have been loosened in readiness - Les Harper?)
  • Finish anything left on the boiler clean out.
  • All boiler plugs and doors to clean: use wire brush on grinding machine in the MPD running shed where boiler washouts are done (use new full face screen & gloves)
  • Left side piston packing and spring to clean off carbon (left on slide bar): try and remove bush (still in gland) carefully.
  • Clean both cylinder and valve front faces and wire brush and grease studs.
  • Remove all carbon from ports on valves & cylinders.
  • Remove all carbon from left piston head.
  • Clean all valve & cylinder covers but take care not to damage the copper ring on cylinder covers.